OpenStudy (anonymous):

gardener harvested 100 kilograms of tomatoes and drove them to the farmer's market on a hot day. When picked the tomatoes were 99% water but they lost some moisture during the drive and were 95% water when delivered. What was their weight when they arrived?

OpenStudy (anonymous):

95/99 = x/100, solve for x x = 95 * 100 / 99 x = 95.96

OpenStudy (anonymous):

hold the phone they were originally 99% water so 1 kg tomatoes, 99 kg water.

OpenStudy (anonymous):

when they arrive there is still 1 kg tomato, the rest water. so the question is, 1 kg is 5% of what number? in other words solve \[1=.05x\] and get \[x=\frac{1}{.05}=20\] there are only 20 kg on arrival

OpenStudy (anonymous):

not the answer you were expecting i am sure, but it is correct